# Elasticity of a Freely Jointed Chain with Quenched Disorder

### v1

@@ 1 Introduction
Quite often, the macroscopic behavior of a system is affected by disorder, and the type of disorder plays an important role. Thus, the difference between the types of disorder associated with certain degrees of freedom has been of interest to various systems, including macromolecules, both synthetic and biopolymers. The most well-known and also well-studied problem is that of polypeptides undergoing helix-coil transition. In the helix state, the monomers form a rotating pattern induced by hydrogen bonds that increase the bending stiffness, whereas, in the coil state, the monomers are oriented randomly, which we can be viewed - for zero or small tension - as a Gaussian chain.
In [1], the mean-field version of the free energy density and the force-extension relation were analyzed, for annealed disorder, treating the helix state as a longer rod (infinite bending stiffness) and the coil state as a freely jointed chain. The quenched-disorder case was also considered, where the freely jointed chain consists of monomers (rods) with two different lengths. The difference in the monomer length is the quenched disorder. [2] treats the annealed version of the same model exactly and points out that there are stress-induced helix-coil crossovers for finite values of the cooperativity parameter, where the transition to the helix state stiffens the chain. A genuine phase transition appears only in the fully cooperative case. [3] considered semi-flexible A-B block copolymers under tension, also treating both types of disorder. The semi-flexible blocks are modeled as wormlike-chain (WLC) segments with different persistence lengths. The approach was based on the Markov process method introduced be Fredrickson, Milner, and Leibler in [4]. The same approach, applied to randomly flexible heteropolymers without tension, was presented in [5]. A numerical study of the conformational and elastic properties of the Kratky-Porod chain with quenched disorder in the bending stiffness was presented in [6].
@@
In a recent work, we have analyzed the stretching and bending elasticity of the freely joined chain with reversible hinges (rFJC) in the Gibbs ensemble [11, 12], which is the annealed scenario of the model we will present in this article. Actually, this model (the annealed version) was originally introduced and analyzed in the Helmholtz (fixed extension) ensemble by Winkler et al. [13, 14]
This model is the simplest model that captures the effect of a random (annealed) bending stiffness. An open hinge acts the same as in the usual freely joined chain (uFJC), and a closed hinge links the two adjacent segments (which are rigid rods), forming a longer rod with infinite bending stiffness. In the thermodynamic limit of the rFJC, it turns out that the force-extension relation in the large force limit becomes the form of $1-{\rm e}^{-f}$, i.e. exponential decay in the differential tensile compliance. This is qualitatively different from the uFJC, where the compliance decays as a power law.
The simplest way to introduce quenched disorder in the local bending stiffness is to make an analogy with the rFJC. This can be accomplished by letting the reversible hinges of the rFJC get quenched. This is the freely joined chain with quenched hinges, and also the paper’s main topic. Then, one can naturally ask the following questions: How will the force-extension relation be different from that of the uFJC or the rFJC? Is there a general framework to deal with quenched disorder on polymers? In this article, we try to gain insight into these questions.
The article is organized as follows. In Section 2, we introduce the model of a freely joined chain with quenched hinges in detail. In Section 3, we discuss a simple case of the model with one random hinge and obtain a recurrence relation of the exact mean free energy. In Section 4, we derive a mathematical form to deal with the thermodynamic limit for quenched-disorder polymers and obtain the force-extension relation for the discrete and continuous models. Finally, in Section 5, we obtain a force-extension law for random copolymers and show an example of how this approach can be implemented.
@@ 2 Model
We start by considering a freely joined chain consisting of $N+1$ chain segments, each of which is a one-dimensional massless rigid rod of length $b$, where the segments are connected by hinges $N$. The total contour length is $L=(N+1)b$. We consider the chain to be in three dimensions, allowing for any orientation of the segments along the unit sphere.
@@ 3.1 Toy System
Instead of heading to the thermodynamic limit immediately, it is helpful to consider a toy system consisting of a few hinges to get a grasp of the situation. However, because of the finite size of the system, the free energy and the extension of the toy system are not self-averaging. Thus, it may be risky to examine the behavior of such a toy system, but it is still useful to gain insight into the effect of quenched disorder. In this subsection, we consider a system with $N=1$, which is sufficient for our purpose.
@@
where we have set $b=1$,$\,\beta=1$ for simplicity. Notice that if $p=0$, it reduces to the free energy of the uFJC as we have mentioned at (4). The force-extension relation in the Gibbs ensemble is given by
@@ 3.2 Recurrence Relation
Using the recurrence relation obtained in (11), it is possible to calculate the exact stretching response for a finite number of hinges numerically. The extension of $N=17$ qFJC is plotted in figure 2. As expected from Section 3.1, we observe that the qFJC is “slower” than the rFJC under the same value of $p$. Note that we have used $p={\rm e}^{\epsilon}/({\rm e}^{\epsilon}+4\pi)$ for the rFJC ($\epsilon$ being the activation energy for opening a reversible hinge), which is the closing probability when $f=0$. Of course, both the rFJC and the qFJC have larger extension than the uFJC for the same force, due to the presence of longer rods.
@@ 4.2 Discrete qFJC
Let us now focus on $\mathcal{P}(k)$. To form a length a block of length $kb$, we should have $k-1$ hinges consecutively closed, and the $k\,$th hinge should be open. Assuming that the opening-closing of each hinge are independent events, it leads to the weighting function
@@
By analogy to the force-extension relation of the uFJC at the strong-force limit, we see that the response expressed by (19) is similar to that of a uFJC with monomer (elementary rod) length equal to $b/(1-p)$.
We see that the resulting force-extension relation qualitatively behaves like a Langevin function, in the sense that the linear response is proportional to the mean cluster size, and it approaches 1 with an inverse-force decay. On the other hand, the rFJC approaches 1 with an exponential decay, which implies that the extension is faster than in the qFJC.
@@ 4.3 Continuous qFJC
One generalization of the quenched Freely Jointed Chain is considering the continuum limit, which is also discussed in [12]. A specific realization of such a chain is drawn in figure 3. For the continuum limit, we have the weighting function as $\mathcal{P}(l)={\rm e}^{-l/L_{p}}/L_{p}$, where $l$ is the block length and $L_{p}$ is the persistence length. Note that $\mathcal{F}(l)$ is given by (14), with changing the variable as $l=kb$. Then, the force-extension relation reads
@@
Surprisingly, this non-trivial integral can be cast into a relatively simple form by some manipulations and using a result of [18], we obtain
@@
It is also interesting to examine the behavior at the limits of small and large force. For the former, where $\alpha/2\rightarrow\infty$, by using the expansion of the gamma function, we obtain
@@
Note that $L_{p}=b/(1-p)$ in the continuum limit. [12] The result above can be confirmed by the linear response theory, which states that $\langle z\rangle\propto\langle R^{2}\rangle f$ where $\langle R^{2}\rangle\approx 2L_{p}L\;(L\gg L_{p})$ is the mean squared end-to-end distance. [12] Note also that
@@
when $p\rightarrow 1$[19] This implies that the result of the continuum limit becomes identical with that of the the discrete chain at the limit of $p\rightarrow 1$. This is somehow expected from the fact that the continuum limit implies $N\rightarrow\infty$ and $b\rightarrow 0$ but keeps $L_{p}=b/(1-p)$ constant.
@@
where $\zeta(z)$ is the Riemman zeta function. Notice that the leading order of (24) becomes identical with (19).
@@ 5.1 Copolymers with Non-Interacting Blocks
The expression for the free energy density in the thermodynamic limit (13) is for homopolymers with non-interacting blocks. The quenched disorder is related to the size of the blocks that follows a certain random distribution. We can generalize our analysis by considering two or more types on non-interacting random blocks, each following a different distribution.
@@
In this case, the number of monomers in the chain will be $N=(S_{1}+S_{2})N_{B}$, where $S_{1}$ and $S_{2}$ are the mean sizes of type 1 and type 2 blocks, respectively. This leads to the mean free energy density
@@
Now consider the force-extension relation of the polymer. In the Gibbs ensemble, the extension reads
@@ 6 Conclusion and Discussion
In this article, we have introduced and analyzed the elastic behavior of the qFJC, a freely-jointed chain with quenched disorder in the closed-open state of its hinges. The qFJC captures the quenched disorder in the local bending stiffness of polymers. Analyzing a toy system with a few random hinges, we observed that the response is the average of the response of the realizations of disorder, as in other systems with quenched disorder. Note that there is a one-to-one correspondence between the possible configurations of the qFJC and the ordered partitioning of integers. This can be understood if we consider the hinges as the possible space to insert partitions. Then, the number of segments between the partitions becomes the size of the integer.
@@
The recurrence relation that we derived for the mean free energy allows us to calculate the exact force-extension relation numerically, although it becomes inefficient for a large value of random hinges $N$. As we expected, it turns out that the force-extension relation is still quantitatively similar to a Langevin function, with a linear response at small forces and an inverse-force limit at strong stretching. If we compare with the rFJC, the rFJC is always faster (more extended) than the qFJC for any value of $f$.
@@
We have derived the exact mean free energy density in the thermodynamic limit, making an analogy with the percolation theory. The result is expressed as an infinite series, and in most cases, it turns out that it is hard to use it directly. However, we obtain simple expressions at the linear-response and the strong-stretching limits. The former is the same as in the rFJC. The latter is qualitatively different from the rFJC and behaves as a uFJC with an effective monomer length.
@@
The mathematical form that we have derived for the discrete qFJC is general and can also be applied to the continuum limit. Remarkably, the continuum limit of the qFJC yields a closed expression for the force-extension relation, involving the trigamma function.
@@
As expected, the response function for the qFJC is always slower than that of the rFJC for both the discrete and the continuous case also in the thermodynamic limit. In the reversible chain (rFJC), there is a crossover to the state with larger bending stiffness (all hinges getting closed) when the force increases, resulting in the formation of a longer rod with a faster response. [12] In contrast, since the disorder architecture is quenched, we do not have such a crossover. This is critical for understanding the slower response for the qFJC.
@@
Noticing that the derivation of (13) is based on the assumption of non-interacting blocks, we generalized to copolymers consisting of different types of random blocks, each having a different length distribution. We have shown that the resulting force-extension relation becomes the average of the extension function over all types. When we extended our analysis to model rod-coil multiblock copolymers, it turned out that the length distribution of the rod affects the force-extension relation, especially in the small-force regime, compared to the case of monodisperse rods.

### v2

@@ 1 Introduction
Quite often, the macroscopic behavior of a system is affected by disorder, and the type of disorder plays an important role. Thus, the difference between the types of disorder associated with certain degrees of freedom has been of interest to various systems, including macromolecules, both synthetic and biopolymers. The most well-known and also well-studied problem is that of polypeptides that undergo helix-coil transition. In the helix state, the monomers form a rotating pattern induced by hydrogen bonds that increase the bending stiffness, whereas, in the coil state, the monomers are oriented randomly, which can be viewed - for zero or small tension - as a Gaussian chain.
In [1], the mean-field version of the free energy density and the force-extension relation were analyzed, for annealed disorder, treating the helix state as a longer rod (infinite bending stiffness) and the coil state as a freely jointed chain. The quenched-disorder case was also considered, where the freely jointed chain consists of monomers (rods) with two different lengths. The difference in the monomer length is the quenched disorder. [2] treats the annealed version of the same model exactly and points out that there are stress-induced helix-coil crossovers for finite values of the cooperativity parameter, where the transition to the helix state stiffens the chain. A genuine phase transition appears only in the fully cooperative case. [3] considered semi-flexible A-B block copolymers under tension, also treating both types of disorder. The semi-flexible blocks are modeled as wormlike-chain (WLC) segments with different persistence lengths. The approach was based on the Markov process method introduced by Fredrickson, Milner, and Leibler in [4]. The same approach, applied to randomly flexible heteropolymers without tension, was presented in [5]. A numerical study of the conformational and elastic properties of the Kratky-Porod chain with quenched disorder in the bending stiffness was presented in [6].
@@
In a recent work, we have analyzed the stretching and bending elasticity of the freely jointed chain with reversible hinges (rFJC) in the Gibbs ensemble [11, 12], which is the annealed scenario of the model we will present in this article. Actually, this model (the annealed version) was originally introduced and analyzed in the Helmholtz (fixed extension) ensemble by Winkler et al. [13, 14]
This model is the simplest model that captures the effect of a random (annealed) bending stiffness. An open hinge acts the same as in the usual freely jointed chain (uFJC), and a closed hinge links the two adjacent segments (which are rigid rods), forming a longer rod with infinite bending stiffness. In the thermodynamic limit of the rFJC, it turns out that the force-extension relation in the large force limit becomes the form of $1-{\rm e}^{-f}$, i.e. exponential decay in the differential tensile compliance. This is qualitatively different from the uFJC, where the compliance decays as a power law.
The simplest way to introduce quenched disorder in the local bending stiffness is to make an analogy with the rFJC. This can be accomplished by letting the reversible hinges of the rFJC get quenched. This is the main topic of the paper, which is the freely jointed chain with quenched hinges. Then, one can ask the following questions: How will the force-extension relation differ from the uFJC or the rFJC? Is there a general framework to deal with quenched disorder on polymers? In this article, we try to gain insight into these questions.
The article is organized as follows. In Section 2, we introduce the model of a freely jointed chain with quenched hinges in detail. In Section 3, we discuss a simple case of the model with one random hinge and obtain a recurrence relation for the exact mean free energy. In Section 4, we derive a mathematical form to deal with the thermodynamic limit for quenched-disorder polymers and obtain the force-extension relation for the discrete and continuous models. Finally, in Section 5, we obtain a force-extension law for random copolymers and show an example of how this approach can be implemented.
@@ 2 Model
We start by considering a freely jointed chain consisting of $N+1$ chain segments, each of which is a one-dimensional massless rigid rod of length $b$, where the segments are connected by hinges $N$. The total contour length is $L=(N+1)b$. We consider the chain to be in three dimensions, allowing for any orientation of the segments along the unit sphere.
@@ 3.1 Toy System
Instead of heading to the thermodynamic limit immediately, it is helpful to consider a toy system consisting of a few hinges to get a grasp of the situation. However, because of the finite size of the system, the free energy and the extension of the toy system are not self-averaging. (The problem of self-averaging in finite systems is discussed later, in Section 3.3.) Thus, it may be risky to examine the behavior of such a system, but it is still useful to gain insight into the effect of quenched disorder. In this subsection, we consider a system with $N=1$, which is sufficient for our purpose.
@@
where we have set $b=1$,$\,\beta=1$ for simplicity. Notice that if $p=0$, it reduces to the free energy of the uFJC as we have mentioned in (4). The force-extension relation in the Gibbs ensemble is given by
@@ 3.2 Recurrence Relation
Using the recurrence relation obtained in (11), it is possible to calculate the exact stretching response for a finite number of hinges numerically. The extension of $N=17$ qFJC is plotted in figure 2. As expected from Section 3.1, we observe that the qFJC is less extended than the rFJC under the same value of $p$. Note that we have used $p={\rm e}^{\epsilon}/({\rm e}^{\epsilon}+4\pi)$ for the rFJC ($\epsilon$ being the activation energy for opening a reversible hinge), which is the closing probability when $f=0$. Of course, both the rFJC and the qFJC have larger extensions than the uFJC for the same force, due to the presence of longer rods.
@@
3.3 Self-Averaging of Finite Systems
Self-averaging plays an important role in describing a system with quenched disorder by its mean value. For random polymers, self-averaging of a certain property implies independence of that property on the random sequence (realization of disorder). In the thermodynamic limit, most of the physical quantities of a system without long-range correlations, such as the qFJC, are self-averaging due to the central limit theorem. However, for finite systems or near criticality, the problem of self-averaging becomes nontrivial. Self-averaging in random copolymers modeled as self-avoiding random walks has been studied by Whittington and collaborators. [17, 18, 19] They have derived a bound on the extent of self-averaging as a function of the polymer length. In this subsection, we estimate a minimal length of the qFJC beyond which self-averaging is a good approximation.
In principle, to be (at least approximately) self-averaging, the system should be large enough to realize most of the possible states that the system can have. Thus, it would be natural to try to find a minimal length that is sufficient to realize most of the possibilities. In our model, this length will be characterized by the closing probability of the hinge $p$.
Let us denote this minimal length as $L_{m}$ and define it as the contour length over which the polymer is likely to have at least two blocks, given that the probability of having a closed hinge is $p$. This definition can be written as follows.
${\rm Prob}({\rm at\;least\;two\;blocks})>\frac{1}{2}\;\;{\rm or}\;\;{\rm Prob}({\rm one\;block})<\frac{1}{2}.$ (12)
If a system has $N$ hinges, the statement above can be written as $p^{N}<1/2$, which is very simple. Thus, we conclude that the self-averaging is a good approximation if the number of hinges $N$ satisfies
$Nb>L_{m}=-\frac{\ln 2}{\ln p}b.$ (13)
We point out that this is just a rough estimate that illustrates the dependence of the minimal length on $p$. One may consider a different requirement for the number of blocks, for example, “at least X blocks”. However, that would yield a transcendental equation.
@@ 4.2 Discrete qFJC
Let us now focus on $\mathcal{P}(k)$. To form a length a block of length $kb$, we should have $k-1$ hinges consecutively closed, and the $k\,$th hinge should be open. Assuming that the opening-closing of each hinge as independent events, it leads to the weighting function
@@
Before we proceed, note that the weighting function $\mathcal{P}(k)$ has a similar form (except for a multiplicative constant) to the orientational correlation function for the discrete chain, which is calculated in 12. This is somehow expected. If we define $\vec{u}(i)$ as the tangent vector of the $i$th segment, $\vec{u}(i)\cdot\vec{u}(j)$ can have two values: 00 or $1$. For the latter, the hinges between the $i$th segment and the $j$th segment should be consecutively closed, which corresponds to the case where the size $|i-j|+1$ block has formed. The former corresponds to the other cases, which give zero contribution to the orientational correlation function. Thus, we have $\langle\vec{u}(i)\cdot\vec{u}(j)\rangle\sim\mathcal{P}(|i-j|)$.
@@
By analogy to the force-extension relation of the uFJC at the strong-force limit, we see that the response expressed by (21) is similar to that of a uFJC with monomer (elementary rod) length equal to $b/(1-p)$. Notice that the response in the two regimes ((20) and (21)) can be expressed by that of a uFJC with two different sizes of effective monomers, $(1+p)b/(1-p)$ and $b/(1-p)$, respectively.
We see that the resulting force-extension relation qualitatively behaves similar to a Langevin function, in the sense that the linear response is proportional to the mean cluster size. However, it approaches full extension with an inverse-force decay with an effective monomer size which is different from the mean cluster size. Thus, it is not exactly true that the tensile elasticity of a qFJC reduces to that of a uFJC with an effective monomer size.
We can prove that the extension of the rFJC is larger than in the qFJC by arguing that both force-extension relations are monotonic functions of the force. Since the extension of the rFJC approaches 1 with an exponential decay and both the qFJC and the rFJC have the same linear response, that automatically implies that the extension is larger than the qFJC under the same conditions.
@@ 4.3 Continuous qFJC
One generalization of the quenched Freely Jointed Chain is considering the continuum limit, which is also discussed in [12]. A specific realization of such a chain is drawn in figure 3. For the continuum limit, we have the weighting function as $\mathcal{P}(l)={\rm e}^{-l/L_{p}}/L_{p}$, where $l$ is the block length and $L_{p}$ is the persistence length. As we discussed in Section 4.2, this weighting function can be justified by considering the orientational correlation function for the continuum limit, which has the form of $\exp(-L/L_{p})$.
The free energy of a length $l$ block, $\mathcal{F}(l)$ is given by (16), with changing the variable as $l=kb$. Then, the force-extension relation reads
@@
Surprisingly, this non-trivial integral can be cast into a relatively simple form by some manipulations, and using a result of [21], we obtain
@@
It is also interesting to examine the behavior at the limits of small and large forces. For the former, where $\alpha/2\rightarrow\infty$, by using the expansion of the gamma function, we obtain
@@
Note that $L_{p}=b/(1-p)$ in the continuum limit. [12] The result above can be confirmed by the linear response theory, which states that $\langle z\rangle\propto\langle z^{2}\rangle f$ where $\langle z^{2}\rangle\approx 2L_{p}L/3\;(L\gg L_{p})$ is the mean squared end-to-end distance. [12] Note also that
@@
when $p\rightarrow 1$[22] This implies that the result of the continuum limit becomes identical to that of the discrete chain at the limit of $p\rightarrow 1$. This is somehow expected from the fact that the continuum limit implies $N\rightarrow\infty$ and $b\rightarrow 0$ but keeps $L_{p}=b/(1-p)$ constant.
@@
where $\zeta(z)$ is the Riemann zeta function. Note that, also in this case, it cannot be viewed as an extension of a uFJC with an effective monomer length of size $L_{p}$, due to the continuous nature of this model. Moreover, even if we try to express it as an uFJC, those two regimes have different effective monomer lengths, $2L_{p}$ (the Kuhn length of a flexible chain) and $L_{p}$.
@@ 5.1 Copolymers with Non-Interacting Blocks
The expression for the free energy density in the thermodynamic limit (15) is for homopolymers with non-interacting blocks. The quenched disorder is related to the size of the blocks that follow a certain random distribution. We can generalize our analysis by considering two or more types on non-interacting random blocks, each following a different distribution.
@@
In this case, the number of monomers in the chain will be $N=(S_{1}+S_{2})N_{B}$, where $S_{1}$ and $S_{2}$ are the mean sizes of blocks of type 1 and type 2, respectively. This leads to the mean free energy density
@@
Now we consider the force-extension relation of the polymer. In the Gibbs ensemble, the extension reads
@@ 6 Conclusion and Discussion
In this article, we have introduced and analyzed the elastic behavior of the qFJC, a freely-jointed chain with quenched disorder in the closed-open state of its hinges. The qFJC captures the quenched disorder in the local bending stiffness of the polymer. We calculated the response by averaging the response over all realizations of the disorder, as in other systems with quenched disorder. Note that there is a one-to-one correspondence between the possible configurations of the qFJC and the ordered partitioning of integers. This can be understood if we consider the hinges as the possible sites of the partitions. Then, the number of segments between the partitions becomes the size of the integer.
@@
The recurrence relation that we derived for the mean free energy allows us to calculate the exact force-extension relation numerically, although it becomes inefficient for a large value of random hinges $N$. We expect our results to be approximately self-averaging, if the contour length is longer than a certain minimal length $L_{m}$ that depends on the closing probability of a hinge $p$. As we expected, it turns out that the force-extension relation is still qualitatively similar to a Langevin function, with a linear response at small forces and an inverse-force limit at strong stretching. Compared to the rFJC, the qFJC is always less extended than the rFJC for any value of the stretching force $f$.
@@
We have derived the exact mean free energy density in the thermodynamic limit, making an analogy with the percolation theory. The result is expressed as an infinite series, and in most cases, it turns out that it is hard to use it directly. However, we obtain simple expressions at the linear-response and the strong-stretching limits. The former is the same as in the rFJC. The latter is qualitatively different from the rFJC and behaves as a uFJC with an effective monomer length. We should point out that both (19) and (22) could be ”derived” directly, without any detour through free energy calculations, by considering a weighted superposition of Langevin functions.
@@
The mathematical form we have derived for the discrete qFJC is general and can also be applied to the continuum limit. Remarkably, the continuum limit of the qFJC yields a closed expression for the force-extension relation, involving the trigamma function.
@@
As expected, the projection of the mean end-to-end distance on the axis of the tensile force for the qFJC is always less extended than that of the rFJC for both the discrete and the continuous case, also in the thermodynamic limit. In the reversible chain (rFJC), there is a crossover to the state with larger bending stiffness (all reversible hinges getting closed) when the force increases, resulting in the formation of a longer rod with a larger response. [12, 11] In contrast, since the disorder architecture is quenched, we do not have such a crossover. This is crucial for understanding the smaller response for the qFJC.
If we want to experimentally determine the type of disorder for this kind of system, we must focus on the large-force regime. Although the initial linear response of the qFJC and the rFJC is the same, the two models behave qualitatively differently in the large-force regime. For both discrete and continuous qFJC, the chain extension approaches 1 (full extension) as $1-A/f$, whereas the rFJC approaches 1 as $1-B\exp(-kf)$ (where $A,\,B$, and $k$ are constants) in the thermodynamic limit.
@@
Noticing that the derivation of (15) is based on the assumption of non-interacting blocks, we generalized to copolymers consisting of different types of random blocks, each having a different length distribution. We have shown that the resulting force-extension relation becomes the average of the extension function over all types of random blocks. When we extended our analysis to model rod-coil multiblock copolymers, it turned out that the length distribution of the rod affects the force-extension relation, especially in the small-force regime, compared to the case of monodisperse rods.
The rFJC and qFJC can be viewed as minimal models of polymers where the local bending stiffness fluctuates along the polymer backbone. For example, some DNA-binding proteins may bind/unbind along the single- or double-stranded DNA and increase its local bending stiffness. [23, 24, 25] If the typical time scale of the binding/unbinding process is smaller than the relaxation time of the DNA, the rFJC will be a more appropriate model. In the opposite case, the qFJC may be more useful. In DNA(or mRNA) that undergoes translation(transcription), the time scale of the translation(transcription) is separated from the relaxation time of the polymer. [26] The presence of a ribosome(polymerase) on the polymer backbone increases its local bending stiffness. For a time scale smaller than the typical transport time of such a process, the qFJC may be a useful model to describe the tensile elasticity.